OpenStudy (anonymous):

Write an expression for the following real-world situation. Katie is moving to a new house. Small boxes hold 15 pounds each and medium boxes hold 25 pounds each. Write an expression for the total weight in the moving van if Katie uses "s" small boxes and "m" medium boxes.

OpenStudy (anonymous):

1 Small box holds 15 pounds Therefore s small boxes will hold 15*s pounds Similarly For large boxes Then just sum up .. Now can u do it ??

OpenStudy (anonymous):

nope this is hard. algebra is not my thing. :(

OpenStudy (anonymous):

See it like this ... If u have one 1$ Then its total value will be 1$ only But if u have ten 1$ then total value of that will be ?? Can u give the answer

OpenStudy (anonymous):

10 dollars

OpenStudy (jhannybean):

Match up the "s" and "m" with the weight each type of box holds. "s"mall boxes hold 15, "m"edium boxes hold 25.

OpenStudy (anonymous):

Yes thats exactly u have to do here Now if u have 5 10$ then & 10 1$ then total value of money u have ???

OpenStudy (anonymous):

Try this one

OpenStudy (dan815):

have u ever put icecubes in your milk

OpenStudy (anonymous):

no I /havent But This Is Math

OpenStudy (dan815):

lol

OpenStudy (anonymous):

50+10=60

OpenStudy (anonymous):

U had done correctly above Now in your query 1 Small box holds 15 pounds Therefore "s" small boxes will hold 15*s pounds 1 Large box holds 25 pounds Therefore "m" large boxes will hold 25*m pounds Total weight = 15*s+25*m

OpenStudy (anonymous):

See how easy u r getting the things ... All the best . Practice more & u"ll be a champ :)

OpenStudy (anonymous):

so 15s + 25m= total weight

OpenStudy (anonymous):

I by mistake has written large box in place of medium box . Yes . U got that See

OpenStudy (anonymous):

How About This. Write a real-world situation of your own and a VARIABLE expression to go with it.

OpenStudy (dan815):

just make it up

OpenStudy (dan815):

write something with numbers and replace those numbers with x and ys

OpenStudy (anonymous):

yeah it's kind of hard.

OpenStudy (anonymous):

@dan815 : will help u with that :)

OpenStudy (dan815):

I asked only1 x questions, but she gave me n answers i dont understand what she was thinking maybe her z number of friends are teaching her n number of wrong things

OpenStudy (dan815):

:P

OpenStudy (dan815):

There are x number of boxes in the garage but each box has a volume of n m^3 and the volume of the garage is y m^3 how many boxes could fit in ther